# Weighted Birkhoff Averages Accelerate Data-Driven Methods

### v1

@@ 1 Introduction
hoping that as $N$ grows, the average approaches an expectation. This idea lies at the heart of data sampling assumptions for diffusion forecasting [giannakis2019data, berry2015nonparametric, thiede2019galerkin, giannakis2020extraction, berry2015nonparametric2], Koopman operator analysis [korda2018convergence, bramburger2024auxiliary], Dynamic Mode Decomposition [korda2018convergence, colbrook2023beyond, bramburger2024auxiliary], and other approaches such as system identification [russo2024convergence] and power spectra [colbrook2024rigorous, colbrook2025rigged].
However, the classical Birkhoff average converges painfully slowly. In typical cases, the error decays only like $1/N$ [kachurovskii1996rate], and for certain $g$, it can be arbitrarily slower [krengel1978speed]. Long trajectories are therefore needed to obtain reliable results. These slow rates set the pace for data-driven methods built upon them, dictating how much data is required for accurate approximation.
The remedy is simple but powerful: replace the uniform average with a weighted one [tong2025weighted, duignan2023distinguishing, das2018super, das2017quantitative, tong2024exponential]. By tapering the ends of the sum, edge effects are suppressed and convergence can improve dramatically, sometimes to super-polynomial or even exponential rates. Weighted ergodic averages thus provide sharper estimates from the same data and improve a wide range of spectral and operator-learning methods.
@@ 2 Weighted Birkhoff averages
Consider a measure space $(\mathcal{X},\mathcal{A},\mu)$ and a measure-preserving map $f:\mathcal{X}\to\mathcal{X}$. The measure $\mu$ is said to be ergodic if every set $S\in\mathcal{A}$ satisfying $f^{-1}(S)=S$ has measure $\mu(S)\in\{0,1\}$. Birkhoff’s ergodic theorem [brin2002introduction, Theorem 4.5.5] then states that if $\mu$ is ergodic and $g\in L^{1}(\mathcal{X},\mu)$, the time average $B_{N}(g)(x)$ presented in (1.1) converges for $\mu$-almost every $x\in\mathcal{X}$ to the space average
@@
The weight $w$ damps the ends of the sum, giving greater emphasis to terms near the centre $n\sim N/2$, while the uniform average $B_{N}$ treats all terms equally. If $\mu$ is ergodic, then for any $w\in\mathcal{W}$ and $g\in L^{1}(\mathcal{X},\mu)$ we have [duignan2023distinguishing, Proposition 2]
@@
As summarised in [tong2025weighted], the convergence rate can be greatly improved beyond the unweighted $\mathcal{O}(1/N)$ rate. If $f$ is periodic, the rate of convergence in (2.4) is exponential. If $f$ is quasiperiodic and $g\in C^{\infty}$, the rate is $\mathcal{O}(N^{-m})$ for any $m>0$; if $g$ is analytic, it again becomes exponential. For chaotic dynamics, no general analytic rate is known, but empirical evidence (including that below) suggests that the weighted and unweighted averages share the same $\mathcal{O}(1/N)$ behaviour.
Figure 1: Dynamics (top) and averages (bottom) of the driven logistic map (2.5). Weighted (red) and unweighted (blue) time averages of the state variable $x$ are presented as a log-log plot of absolute errors against the average with $N=10^{6}$. Parameter values are $\varepsilon=0$ (left, periodic), $\varepsilon=0.01$ (middle, quasiperiodic), and $\varepsilon=0.1$ (right, chaotic).
@@
To illustrate, consider the driven logistic map [negi2000bifurcations]:
@@
The top row of Figure 1 shows the iterates $x_{n}$ for these three regimes, all with $x_{0}=0.25$. The bottom row plots, on a log-log scale, the absolute error between the running average and the value at $N=10^{6}$. The unweighted average converges at roughly the $\mathcal{O}(1/N)$ rate for all cases. For non-chaotic dynamics ($\varepsilon=0,0.01$), the weighted averages converge exponentially to machine precision; the small oscillations between $10^{-14}$ and $10^{-16}$ reflect floating-point noise. When $\varepsilon=0.1$ the map is chaotic, and the weighted and unweighted averages converge at nearly the same rate. Thus, weighting never slows convergence and can dramatically accelerate it whenever the underlying dynamics are not chaotic.
@@ 3 Five Examples of Weighted Methods
We build on several well-established data-driven methods for analysing dynamical datasets. Background and fundamentals can be found in [bramburger2024databook, colbrook_book, brunton2022data]. Here, we show how these methods can be strengthened by incorporating weighted Birkhoff averages. When a specific weight $w\in\mathcal{W}$ is needed for illustration, we use the bump function defined in (2.2). All examples that follow can be reproduced with the freely available code at https://github.com/jbramburger/weighted_methods.
@@ 3.1 Example Method I: Dynamic mode decomposition
Dynamic Mode Decomposition (DMD) is a data-driven technique for extracting structure from sequential data [tu2014dynamic, schmid2010dynamic]; see also the review [colbrook2024multiverse]. Given snapshots $X_{1},X_{2},\dots,X_{N},X_{N+1}\in\mathbb{R}^{d}$, for some $N,d\geq 1$, the goal is to identify a best-fit linear map for $X_{n}\mapsto X_{n+1}$. In its simplest form, DMD seeks a matrix $A\in\mathbb{R}^{d\times d}$ so that $X_{n+1}\approx AX_{n}$ for all $n=1,\dots,N$ by minimizing $\|\mathbb{Y}-A\mathbb{X}\|_{F}$, where $\|\cdot\|_{F}$ is the Frobenius norm and
@@
The optimal solution is the DMD matrix $A=\mathbb{Y}\mathbb{X}^{\dagger}$, where $\dagger$ denotes the Moore–Penrose pseudoinverse. The eigenvectors of $A$ (DMD modes) represent coherent spatial structures that grow, decay, or oscillate linearly in time according to their associated eigenvalues.
This matrix can be written equivalently as
@@
Convergence results for DMD as $N\to\infty$ [korda2018convergence, colbrook2023beyond, bramburger2024auxiliary] rely on the convergence of these averages; the limiting matrix represents a projection of the Koopman operator onto the snapshot coordinates (see the following subsection).
@@
where $W:=\mathrm{diag}({w(0)},{w(1/N)},\dots,{w(1)})\in\mathbb{R}^{N\times N}$. The weighted DMD matrix is
@@
    Arrange data into matrices $\mathbb{X}$ and $\mathbb{Y}$ defined in (3.1) and $W=\mathrm{diag}(w(0),w(1/N),\dots,w(1))$.
@@
To demonstrate wtDMD, we consider the classical example of laminar vortex shedding behind a circular cylinder [taira2020modal]. The cylinder has unit diameter, and the Reynolds number is set to $\mathrm{Re}=100$, beyond the critical value where a supercritical Hopf bifurcation produces periodic vortex shedding [jackson1987finite, zebib1987stability]. At this Reynolds number, the flow settles on a stable limit cycle. We analyse vorticity data on a rectangular computational domain of size $18\times 5$, with no-slip boundary conditions on the cylinder and on the top and bottom walls. A parabolic inflow is imposed at the left boundary and a non-reflecting outflow at the right. Each snapshot of vorticity consists of $158,624$ numerical values.
@@
The middle and right columns of Figure 2 show the relative errors of the DMD and wtDMD matrices (middle) and their eigenvalues (right) for $N=10,20,\ldots,500$. For small datasets ($N\lesssim 100$), both methods perform similarly, but once $N\gtrsim 200$, the wtDMD errors drop by several orders of magnitude. Thus, weighting provides faster convergence without altering the asymptotic limit, an easy gain in accuracy from the same data.
@@ 3.2 Example Method II: Approximating the Koopman operator
Consider the discrete-time stochastic process defined by the random map $x_{n+1}=f(\omega(n),x_{n})$, where $\omega\mapsto f(\omega,\cdot)$ is a random variable from a probability space $(\Omega,\mathcal{F},P)$ to the space of continuous maps from $\mathcal{X}\subseteq\mathbb{R}^{d}$ to itself. The (stochastic) Koopman operator [wanner2022robust, vcrnjaric2020koopman] acts on bounded continuous observables $\phi:\mathcal{X}\to\mathbb{R}$ via
@@
This formulation includes deterministic dynamics $x_{n+1}=f(x_{n})$ as a special case. In that setting, the Koopman operator acts by composition [mezic2021koopman, mauroy2020koopman]:
@@
The Koopman operator provides a linear, deterministic representation of dynamics that may themselves be nonlinear and even stochastic. The price we pay is dimensional: instead of evolving a point in the finite-dimensional state space $\mathcal{X}$, we now evolve observables $\phi:\mathcal{X}\to\mathbb{C}$ in an infinite-dimensional function space. Over the past decade, there has been intense interest in approximating the Koopman operator from data, both for forecasting and for extracting coherent structures; see [colbrook_book, Chap. 11] and the reviews [brunton2022modern, colbrook2024multiverse].
The most widely used data-driven method for approximating the Koopman operator on a finite set of observables is extended dynamic mode decomposition (EDMD) [williams2015data]. Suppose we are given sequential data $\{X_{n}\}_{n=1}^{N+1}\subset\mathcal{X}$ from a (possibly stochastic) process. We select two dictionaries of bounded, continuous observables,
@@
This choice of $K^{(N)}$ minimizes $\|\Phi-\Psi K\|_{F}$ over all $K\in\mathbb{C}^{L\times R}$. As $N\to\infty$, the matrix $K^{(N)}$ represents the operator $\mathcal{P}_{L}\mathcal{K}\mathcal{P}_{R}$ [korda2018convergence, colbrook2023beyond, bramburger2024auxiliary]. When $\vec{\phi}$ and $\vec{\psi}$ are just the state coordinates, EDMD reduces (up to transpose) to DMD. Standard EDMD typically takes $\vec{\phi}=\vec{\psi}$ to obtain a square matrix for spectral analysis, but allowing $L\neq R$ can be advantageous; for example, in combining EDMD with semidefinite programming to infer Lyapunov functions and invariant measures directly from data [bramburger2024auxiliary, bramburger2024data].
@@
which exposes the Birkhoff averages that control convergence:
@@
    Compute the matrices $\Psi$ and $\Phi$ defined in (3.6) and $W=\mathrm{diag}(w(0),w(1/N),\dots,w(1))$.
@@
Figure 3: Averaged relative norm errors between the EDMD (blue) and wtEDMD (red) over a range of snapshots $N$ and the benchmark results using $N=10^{7}$ snapshots. All results use data from the standard map (3.8) with: (a) $\lambda=0.5$; (b) $\lambda=0.75$; and (c) $\lambda$ drawn from the uniform distribution on $[0,5]$.
Figure 3 reports the relative Frobenius norm error of the $9\times 9$ matrices $K^{(N)}$ (EDMD) and $K_{w}^{(N)}$(wtEDMD), using the result at $N=10^{7}$ as a benchmark. For each $N$, we average the error over 100 initial conditions $(p_{0},\theta_{0})$ drawn uniformly from the torus. In line with the behaviour seen earlier in Figure 1, weighting offers the greatest advantage in the quasiperiodic regime ($\lambda=0.25$), where convergence is dramatically accelerated. For chaotic and stochastic dynamics, the weighted and unweighted methods converge at essentially the same $\mathcal{O}(1/N)$ rate. Thus, weighting does not degrade performance in difficult regimes, and in favourable regimes it can improve accuracy by orders of magnitude. In applications, replacing EDMD with wtEDMD is therefore a cheap upgrade in accuracy.

### v2

@@ 1 Introduction
hoping that as $N$ grows, the average approaches an expectation. This idea lies at the heart of data sampling assumptions for diffusion forecasting [28, 2, 45, 27, 3], Koopman operator analysis [33, 5], Dynamic Mode Decomposition [33, 18, 5], and other approaches such as system identification [41] and power spectra [17, 16].
However, the classical Birkhoff average converges painfully slowly. In typical cases, the error decays only like $1/N$ [31], and for certain $g$, it can be arbitrarily slower [34]. Long trajectories are therefore needed to obtain reliable results. These slow rates set the pace for data-driven methods built upon them, dictating how much data is required for accurate approximation.
The remedy is simple but powerful: replace the uniform average with a weighted one [47, 26, 25, 24, 46]. By tapering the ends of the sum, edge effects are suppressed and convergence can improve dramatically, sometimes to super-polynomial or even exponential rates. Weighted ergodic averages thus provide sharper estimates from the same data and improve a wide range of spectral and operator-learning methods.
@@ 2 Weighted Birkhoff averages
Consider a measure space $(\mathcal{X},\mathcal{A},\mu)$ and a measure-preserving map $f:\mathcal{X}\to\mathcal{X}$. The measure $\mu$ is said to be ergodic if every set $S\in\mathcal{A}$ satisfying $f^{-1}(S)=S$ has measure $\mu(S)\in\{0,1\}$. Birkhoff’s ergodic theorem [8, Theorem 4.5.5] then states that if $\mu$ is ergodic and $g\in L^{1}(\mathcal{X},\mu)$, the time average $B_{N}(g)(x)$ presented in (1.1) converges for $\mu$-almost every $x\in\mathcal{X}$ to the space average
@@
The weight $w$ damps the ends of the sum, giving greater emphasis to terms near the centre $n\sim N/2$, while the uniform average $B_{N}$ treats all terms equally. If $\mu$ is ergodic, then for any $w\in\mathcal{W}$ and $g\in L^{1}(\mathcal{X},\mu)$ we have [26, Proposition 2]
@@
As summarised in [47], the convergence rate can be greatly improved beyond the unweighted $\mathcal{O}(1/N)$ rate. If $f$ is periodic, the rate of convergence in (2.4) is exponential. If $f$ is quasiperiodic and $g\in C^{\infty}$, the rate is $\mathcal{O}(N^{-m})$ for any $m>0$; if $g$ is analytic, it again becomes exponential. For chaotic dynamics, no general analytic rate is known, but empirical evidence (including that below) suggests that the weighted and unweighted averages share the same $\mathcal{O}(1/N)$ behaviour.
For smooth quasiperiodic dynamics, the slow convergence of Birkhoff averages is mainly a boundary artifact caused by the sharp truncation of the time sum. The average behaves like a low-order quadrature applied to a smooth periodic function, with errors dominated by endpoint effects rather than interior smoothness. Weighted averages replace this abrupt cutoff with a smooth window whose derivatives vanish at the endpoints, suppressing oscillatory Fourier modes by repeated integration by parts. Each non-constant frequency is then damped faster than any algebraic power of the averaging length, while the zero mode (the desired average) is unchanged. In effect, the weighting upgrades the time average to a high-order quadrature rule, yielding super-polynomial convergence for smooth quasiperiodic systems with Diophantine frequencies. This mechanism relies on the regular spectral structure of quasiperiodicity and does not generally extend to chaotic dynamics, where broadband correlations dominate.
Figure 1: Dynamics (top) and averages (bottom) of the driven logistic map (2.5). Weighted (red) and unweighted (blue) time averages of the state variable $x$ are presented as a log-log plot of absolute errors against the average with $N=10^{8}$. Parameter values are $\varepsilon=0$ (left, periodic), $\varepsilon=0.01$ (middle, quasiperiodic), and $\varepsilon=0.1$ (right, chaotic).
@@
To illustrate, consider the driven logistic map [40]:
@@
The top row of Figure 1 shows the iterates $x_{n}$ for these three regimes, all with $x_{0}=0.25$. The bottom row plots, on a log-log scale, the absolute error between the running average and the value at $N=10^{8}$. The unweighted average converges at roughly the $\mathcal{O}(1/N)$ rate for all cases. For non-chaotic dynamics ($\varepsilon=0,0.01$), the weighted averages converge exponentially to machine precision; the small oscillations between $10^{-14}$ and $10^{-16}$ reflect floating-point noise. When $\varepsilon=0.1$ the map is chaotic, and the weighted and unweighted averages converge at nearly the same rate. Thus, weighting never slows convergence and can dramatically accelerate it whenever the underlying dynamics are not chaotic.
@@ 3 Five Examples of Weighted Methods
We build on several well-established data-driven methods for analysing dynamical datasets. Background and fundamentals can be found in [7, 21, 11]. Here, we show how these methods can be strengthened by incorporating weighted Birkhoff averages. When a specific weight $w\in\mathcal{W}$ is needed for illustration, we use the bump function defined in (2.2). All examples that follow can be reproduced with the freely available code at https://github.com/jbramburger/weighted_methods.
@@ 3.1 Example Method I: Dynamic mode decomposition
Dynamic Mode Decomposition (DMD) is a data-driven technique for extracting structure from sequential data [48, 42]; see also the review [19]. Given snapshots $X_{1},X_{2},\dots,X_{N},X_{N+1}\in\mathbb{R}^{d}$, for some $N,d\geq 1$, the goal is to identify a best-fit linear map for $X_{n}\mapsto X_{n+1}$. In its simplest form, DMD seeks a matrix $A\in\mathbb{R}^{d\times d}$ so that $X_{n+1}\approx AX_{n}$ for all $n=1,\dots,N$ by minimizing $\|\mathbb{Y}-A\mathbb{X}\|_{F}$, where $\|\cdot\|_{F}$ is the Frobenius norm and
@@
The optimal solution is the DMD matrix $A=\mathbb{Y}\mathbb{X}^{\dagger}$, where $\dagger$ denotes the Moore–Penrose pseudoinverse. The Moore–Penrose pseudoinverse can be computed from the singular value decomposition of a matrix $\mathbb{X}=U\Sigma V^{\top}$ as $\mathbb{X}^{\dagger}=V\Sigma^{\dagger}U^{\top}$ with the pseudoinverse of the rectangular diagonal matrix $\Sigma$, denoted $\Sigma^{\dagger}$, obtained by taking the reciprocal of each non-zero element of $\Sigma$ and leaving the zeros in place. The eigenvectors of $A$ (DMD modes) represent coherent spatial structures that grow, decay, or oscillate linearly in time according to their associated eigenvalues.
Using the pseudoinverse identity $\mathbb{X}^{\dagger}=\mathbb{X}^{\top}(\mathbb{X}\mathbb{X}^{\top})^{\dagger}$, this matrix can be written equivalently as
@@
Convergence results for DMD as $N\to\infty$ [33, 18, 5] rely on the convergence of these averages; the limiting matrix represents a projection of the Koopman operator onto the snapshot coordinates (see the following subsection).
@@
where $W:=\mathrm{diag}(w(0),w(1/N),\dots,w((N-1)/N)))\in\mathbb{R}^{N\times N}$. The weighted DMD matrix is
@@
    Arrange data into matrices $\mathbb{X}$ and $\mathbb{Y}$ defined in (3.1) and $W=\mathrm{diag}(w(0),w(1/N),\dots,w((N-1)/N))$.
@@
To demonstrate wtDMD, we consider the classical example of laminar vortex shedding behind a circular cylinder [44]. The cylinder has unit diameter, and the Reynolds number is set to $\mathrm{Re}=100$, beyond the critical value where a supercritical Hopf bifurcation produces periodic vortex shedding [30, 51]. At this Reynolds number, the flow settles on a stable limit cycle. We analyse vorticity data on a rectangular computational domain of size $18\times 5$, with no-slip boundary conditions on the cylinder and on the top and bottom walls. A parabolic inflow is imposed at the left boundary and a non-reflecting outflow at the right. Each snapshot of vorticity consists of $158,624$ numerical values.
@@
The middle and right columns of Figure 2 show the relative errors of the DMD and wtDMD matrices (middle) and their eigenvalues (right) for $N=10,20,\ldots,500$, measured against the benchmark $N=1000$. To compute eigenvalue errors, the $r$ eigenvalues of each DMD matrix are first sorted in decreasing order according to the real-valued quantity $\lambda\mapsto|1-\lambda|$, and the relative error is then taken with respect to the correspondingly sorted eigenvalues for $N=1000$. For small datasets ($N\lesssim 100$), both methods perform similarly, but once $N\gtrsim 200$, the wtDMD errors drop by several orders of magnitude. Thus, weighting provides faster convergence without altering the asymptotic limit, an easy gain in accuracy from the same data.
@@ 3.2 Example Method II: Approximating the Koopman operator
Consider the discrete-time stochastic process defined by the random map $x_{n+1}=f(\omega(n),x_{n})$, where $\omega\mapsto f(\omega,\cdot)$ is a random variable from a probability space $(\Omega,\mathcal{F},P)$ to the space of continuous maps from $\mathcal{X}\subseteq\mathbb{R}^{d}$ to itself. The (stochastic) Koopman operator [49, 23] acts on bounded continuous observables $\phi:\mathcal{X}\to\mathbb{R}$ via
@@
This formulation includes deterministic dynamics $x_{n+1}=f(x_{n})$ as a special case. In that setting, the Koopman operator acts by composition [39, 37]:
@@
The Koopman operator provides a linear, deterministic representation of dynamics that may themselves be nonlinear and even stochastic. The price we pay is dimensional: instead of evolving a point in the finite-dimensional state space $\mathcal{X}$, we now evolve observables $\phi:\mathcal{X}\to\mathbb{C}$ in an infinite-dimensional function space. Over the past decade, there has been intense interest in approximating the Koopman operator from data, both for forecasting and for extracting coherent structures; see [21, Chap. 11] and the reviews [9, 19].
The most widely used data-driven method for approximating the Koopman operator on a finite set of observables is extended dynamic mode decomposition (EDMD) [50]. Suppose we are given sequential data $\{X_{n}\}_{n=1}^{N+1}\subset\mathcal{X}$ from a (possibly stochastic) process. We select two dictionaries of bounded, continuous observables,
@@
This choice of $K^{(N)}$ minimizes $\|\Phi-\Psi K\|_{F}$ over all $K\in\mathbb{C}^{L\times R}$. As $N\to\infty$, the matrix $K^{(N)}$ represents the operator $\mathcal{P}_{L}\mathcal{K}\mathcal{P}_{R}$ [33, 18, 5]. When $\vec{\phi}$ and $\vec{\psi}$ are just the state coordinates, EDMD reduces (up to transpose) to DMD. Standard EDMD typically takes $\vec{\phi}=\vec{\psi}$ to obtain a square matrix for spectral analysis, but allowing $L\neq R$ can be advantageous; for example, in combining EDMD with semidefinite programming to infer Lyapunov functions and invariant measures directly from data [5, 6].
@@
where the superscript * denotes the conjugate transpose of the matrix, which exposes the Birkhoff averages that control convergence:
@@
    Compute the matrices $\Psi$ and $\Phi$ defined in (3.6) and $W=\mathrm{diag}(w(0),w(1/N),\dots,w((N-1)/N))$.
@@
Figure 3: Averaged relative norm errors between the EDMD (blue) and wtEDMD (red) over a range of snapshots $N$ and the benchmark results using $N=10^{7}$ snapshots. All results use data from the standard map (3.8) with: (a) $\lambda=0.25$; (b) $\lambda=5$; and (c) $\lambda$ drawn from the uniform distribution on $[0,5]$.
We draw a random initial condition $(p_{0},\theta_{0})$ uniformly from the torus and iterate it under the standard map for $n=10^{7}$ steps. From these iterations we take the first $N\leq 10^{6}$ snapshots and form the $9\times 9$ EDMD and wtEDMD matrices $K^{(N)}$ and $K_{w}^{(N)}$. Their relative errors are measured against the $N=10^{7}$ benchmark by
$\frac{\|K^{(N)}-K^{(10^{7})}\|_{F}}{\|K^{(10^{7})}\|_{F}},\quad\frac{\|K_{w}^{(N)}-K_{w}^{(10^{7})}\|_{F}}{\|K_{w}^{(10^{7})}\|_{F}},$
respectively. We repeat this experiment for 100 independently drawn initial conditions and, for each $N$, average the resulting relative errors. The averaged Frobenius-norm errors are shown in Figure 3. In line with the behaviour seen earlier in Figure 1, weighting offers the greatest advantage in the quasiperiodic regime ($\lambda=0.25$), where convergence is dramatically accelerated. For chaotic and stochastic dynamics, the weighted and unweighted methods converge at essentially the same $\mathcal{O}(1/N)$ rate. Thus, weighting does not degrade performance in difficult regimes, and in favourable regimes it can improve accuracy by orders of magnitude. In applications, replacing EDMD with wtEDMD is therefore a cheap upgrade in accuracy.